Assume all temperatures to be exact, and neglect significant figures for small changes in dimension.A diver releases an air bubble of volume 5.0cm3 from a depth of 24m below the surface of a lake, where the temperature is 9.0?C.

What is the volume of the bubble when it reaches just below the surface of the lake, where the temperature is 20?C?
Answer #1

PIVI P2V2 72 T1 ..................(1)

Where p1 = po + ? gh   (at a depth of 24 m below the surfaceof a lake)

          P1 = 101300 Pa + (1000kg/m3)(9.80m/s2)(24 m)

           P1     =336500 Pa

           T1= (273 +9.0) = 282 K

           V1= 5.0 cm3

           P2= Po = 101300 Pa

           T2= (273 +20) = 293 K

          V2 = ?

Plugging these values in Eq.(1), yields

V2 = T2*P1*V1/T1*Po = 293*336500*5/282*101300

V2 = 17.26 cm^3
Answer #2

at a depth 24 m

P1 = Po + rho*g*h = 1.013*10^5 + 1000*9.8*24 = 3.365*10^5 pa

V1 = 5 cm^3

T1 = 9 + 273 = 283 K

below the surface,

P2 = Po = 1.013*10^5

T2 = 20 + 273 = 293 K

V2 = ?

Apply,

P1*V1/T1 = P2*V2/T2

==> V2 = V1*(P1/P2)*(T2/T1)

= 5*(3.365/1.013)*(293/283)

= 17.2 cm^3 <<<<<<<<<<<<<<<<<<--------------Answer